OpenStudy (anonymous):

Dewayne put a container of food at twenty-five degrees Celsius into the freezer until its temperature measured negative eighteen degrees Celsius. How much did the temperature of the food decrease? seven degrees Celsius eighteen degrees Celsius twenty-five degrees Celsius forty-three degrees Celsius

jimthompson5910 (jim_thompson5910):

Subtract the two temperatures 25 - (-18) = ???

OpenStudy (anonymous):

7?

jimthompson5910 (jim_thompson5910):

That would be true IF it was 25 - 18, but it's not

OpenStudy (triciaal):

43

jimthompson5910 (jim_thompson5910):

Remember two negatives make a positive, so 25 - (-18) = 25 + 18

OpenStudy (anonymous):

i think its 43

jimthompson5910 (jim_thompson5910):

that is correct